MICROWAVE ROTATIONAL SPECTRA FOR A DINUCLEAR COMPLEX - MANGANESE RHENIUM DE-CACARBONYL ^{a}

\underline{CHAKREE TANJAROON}, KRISTEN KECK AND STEPHEN KUKOLICH, Department of Chemistry, University of Arizona, Tucson, AZ 85721.

This appears to be the first microwave measurement of the rotational spectrum of a transition metal - dinuclear complex. Rotational transitions in the 4-7 GHz range were measured using a Flygare-Balle type, pulsed-beam Fourier transform spectrometer for manganese-rhenium decacarbonyl [(CO)_{5}-Mn-Re-(CO)_{5}]. Due to the 3Å separation of the metal atoms the B-rotational constants are close to 200 MHz. This resulted in weak transitions above 4 GHz. Quadrupole hyperfine structure due to Rhenium was partially resolved on some of the observed transitions. In the present work transitions have been measured from J=11-12 through J=14-15 and indicate that this complex is a symmetric top. Rhenium quadrupole coupling parameters and rotational constants will be reported.

aSupported by THE NATIONAL SCIENCE FOUNDATION